OpenStudy (anonymous):

Part 1 [2 points] Is the following equation true? Part 2 [6 points] Use complete sentences to explain the properties used in making your decision. 3 log 2 x +1/2 log 2 y – 3 log 2 z = log 2 x^3sqrt y/z^3

OpenStudy (anonymous):

\[3 \log _{2}x +1/2 \log _{2}y-3 \log _{2}z=\log _{2} x ^{3}\sqrt{y}/z^3]

OpenStudy (anonymous):

\(\large 3 \log _{2}x +\frac{1}{2} \log _{2}y-3 \log _{2}z=\log _{2} (x ^{3}\frac{\sqrt{y}}{z^3}) \) that the problem?

OpenStudy (anonymous):

I have this same question! Someone help?